OpenStudy (anonymous):

Integral Calculus: ∫dm/√e^-4m -1

OpenStudy (anonymous):

can someone please tell me how to start? This is the clearer version: \[\int\limits_{}^{} \frac{ dm }{ \sqrt{e ^{-4m}-1}}\]

OpenStudy (anonymous):

After some simplifications, we get \[\int\limits_{}^{} \frac{ e^{2m} dm }{ \sqrt{1-e^{2m}}} \]

OpenStudy (anonymous):

Sorry for asking this, but how did you simplify it?

OpenStudy (anonymous):

Put \[e^{-4m} = \frac{ 1 }{ e^{4m} }\]

OpenStudy (anonymous):

Oh. I see. Thanks!

OpenStudy (anonymous):

HEy sorry it's a mistake there..

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ e^{2m} dm }{ \sqrt{1-e^{4m}} }\]

OpenStudy (anonymous):

Wait. After simplification, does this mean that I can now proceed in using the inverse trigonometric formula for arcsin? \[u ^{2}=e ^{2m}\] but the derivative of u will not be equal to the one on the top. there is still an e. Um what will happen? Thanks!

OpenStudy (anonymous):

put e^2m = u

OpenStudy (anonymous):

then du/2 = e^2m dm

OpenStudy (anonymous):

\[\frac{1}{2} \int\limits_{}^{} \frac{ du }{ \sqrt{1-u^2} }\]

OpenStudy (anonymous):

Thank you very much! :) I'm sorry, I didn't see the corrected version.

OpenStudy (anonymous):

No problem.